OpenStudy (anonymous):

two particles of masses 1g and 4g are moving with equal kinectic energy.what is the ratio of their linear momenta

OpenStudy (anonymous):

not true the velocity is not the same, (1/2)*m1*v1^2=(1/2)*m2*v2^2.... this will give you v1 in terms of v2 and you can then use the ratio above

OpenStudy (anonymous):

v1=[sqrt(m2/m1)]*v2

OpenStudy (anonymous):

oh sorry, I read the question wrong. Yes you are right what I typed above is WRONG!!! I read the velocities were the same. Of course the velcoties are not the same.

OpenStudy (anonymous):

The answer is1:2

OpenStudy (anonymous):

mA=1g mB=4g mB is 4 times larger than mA EcA=(mAvA^2)/2 EcB=(mBvB^2)/2 EcA=EcB (mAvA^2)/2=(4mAvB^2)/2 vA^2=4vB^2 vA=2vB p=linear momentum pA=mAvA pB=mBvB pA/pB=[mAvA]/[mBvB] pA/pB=[mA2vB]/[4mAvB] pA/pB=2/4 pA/pB=1/2

OpenStudy (anonymous):

k=(1/2)mv^2 k=(1/2m)((mv)^2) here mv is linear momenta (mv)^2=2km (m1v1)^2/(m2v2)^2=2km1/2km2=m1/m2=1/4 therefore, (m1v1)/(m2v2)=1/2 hence,ratio of their linear momenta is 1/2

OpenStudy (anonymous):

did u get this?

OpenStudy (anonymous):

plz answer my question..........................